A stock has an expected return of 14 percent, its beta is 1.3, and the expected return on the market is 12 percent. What must the risk-free rate be? (Do not round your intermediate calculations.)
Answer #1

Based on CAPM, Expected return on stock Risk free rate Beta * (Expected market return - Risk free rate)14% = Rf + 1.3 * (12% - Rf)

14% = Rf + 15.6% - 1.3 * Rf

1.6% = 0.3 * Rf

Rf = 5.33%